OpenStudy (anonymous):

HELP! What number would go into the blanks to make a perfect square trinomial? (X^2+10x+_______)=______+6?

OpenStudy (freckles):

\[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\]

Directrix (directrix):

Take half the coefficient of the x-term, square it, and add it to both sides of the equation.